OpenStudy (anonymous):

given that sin theta equals 1/5 and that the terminal side is in quadrant 1, find exact answers for cos(theta+pi/4)

OpenStudy (campbell_st):

well if you draw a diagram  don't you need to find x using pythagoras' theorem.... and subsequently \[\cos(\theta)\] then use the expansion \[\cos(\theta + \frac{\pi}{4}) = \cos(\theta)\cos(\frac{\pi}{4}) - \sin(\theta)\sin(\frac{\pi}{4})\] hope it helps